Required information [The following information applies to the questions displayed below.] Cane Company manufactures two prod2. What is the companys total amount of common fixed expenses? Total common fixed expenses3. Assume that Cane expects to produce and sell 96,000 Alphas during the current year. One of Canes sales representatives ha4. Assume that Cane expects to produce and sell 106,000 Betas during the current year. One of Canes sales representatives ha
Answer #1
1
Alpha Beta
Traceable fixed manufacturing overhead per unit 31 34
X Expected production 125000 125000
Traceable fixed manufacturing overhead 3875000 4250000
2
Total Common fixed expenses 7125000 =125000*(31+26)
3
Variable costs: Alpha
Direct materials 42
Direct labor 35
Variable manufacturing overhead 23
Variable selling expenses 28
Total variable costs 128
Incremental revenue 3744000 =26000*144
Less: Incremental variable costs 3328000 =26000*128
Net change in income 416000
Financial advantage 416000
4
Variable costs: Beta
Direct materials 21
Direct labor 28
Variable manufacturing overhead 21
Variable selling expenses 24
Total variable costs 94
Incremental revenue 296000 =4000*74
Less: Incremental variable costs 376000 =4000*94
Net change in income -80000
Financial (disadvantage) (80000)